A leasing contract calls for an immediate payment of $119,000 and nine subsequent $119,000 semi-annual payments at six-month intervals. What is the PV of these payments if the annual discount rate is 14%?
Answer #1

Given about a leasing contract,

Semi - annual payment PMT = $119000 starting today

Since first payment is today, it is an annuity due.

discount rate = 14% compounded semiannually,

So, semiannual rate r = 14/2 = 7%

total payment N = 10 including today's payment

So, PV of this annuity due is calculated using formula

PV = PMT*(1+r)*(1 - (1+r)^(-N))/r = 119000*1.07*(1 - 1.07^-10)/0.07 = $894312.64

So, PV of these payments = $894312.64